OpenStudy (celinegirl):

Which statement is true about the product square root of 2(3square root of 2 + square root of 18)? It is rational and equal to 12. It is rational and equal to 18. It is irrational and equal to 6 + 2square root of 18. It is irrational and equal to 3 +square root of 18.

OpenStudy (anonymous):

can you post a screenshot or attachment? not really clear from the words

OpenStudy (rccb):

@ANIMALOVER02

OpenStudy (animalover02):

ok give me a sec to figure this out..

OpenStudy (animalover02):

its the third one/please fan and metal

OpenStudy (celinegirl):

Yes I can post a snapshot

OpenStudy (celinegirl):

@ANIMALOVER02 really? I was thinking it was the first one

OpenStudy (animalover02):

if ya want to get it wrong go with your answer but im amazing at math and english

OpenStudy (celinegirl):

@ANIMALOVER02 can you explain?

OpenStudy (anonymous):

yeah it's 1st one

OpenStudy (jchick):

\[\sqrt2(3 \sqrt{2}+18)\]

OpenStudy (jchick):

Right?

OpenStudy (jchick):

a(b+c)=ab+ac

OpenStudy (jchick):

Here it is expanded as much as possible.

OpenStudy (jchick):

So now you know your answer

OpenStudy (celinegirl):

@jchick so it is the third one?

OpenStudy (jchick):

Yes you are correct!